OpenStudy (anonymous):

A soup can has a height of 4.5 inches and a diameter of 3.5 inches. Which of the following is the closest to the volume of the soup can? A. 43 cubic inches B. 173 cubic inches C. 56 cubic inches D. 25 cubic inches @IGREEN I think its A but im not sure

OpenStudy (igreen):

A is correct. \(\sf V = \pi r^2 h\) \(\sf V = (3.14)(1.75^2)(4.5)\) \(sf V = (3.14)(3.0625)(4.5)\) \(\sf V \approx 43\)

OpenStudy (igreen):

It's a cylinder, right?

OpenStudy (anonymous):

yes and thank you @igreen

OpenStudy (igreen):

Thank yourself, you got it right \(\huge\ddot\smile\)